# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v5

@@ V.3 Second Phase: Goodness-of-fit Test
VI Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
@@
The statistics discussed in Eq. (12) of Section IV consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section VII.
VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
@@
Currently available photon arrival time-based QRNGs quRNG ; TROPOS use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P(i;1st|n)$ $=\frac{\sum_{l=1}^{n}P(0,\frac{(i-1)T}{N})P(l,\frac{T}{N})P(n-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum_{l=1}^{n}e^{-\mu T\frac{i-1}{N}}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-l}\left(1-\frac{i}{N}\right)^{n-l}}{(n-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=\sum_{l=1}^{n}\binom{n}{l}\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-l}$
$=\left(1-\frac{i-1}{N}\right)^{n}-\left(1-\frac{i}{N}\right)^{n}.$ (26)
This was also observed by Nie et al. (10.1063/1.4863224, , Eq. (2)). However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
@@
VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
@@
$=\frac{\sum\limits_{k=0}^{n-1}\frac{e^{-\mu T\frac{i-1}{N}}\left(\mu T\frac{i-1}{N}\right)^{k}}{k!}(1-d)^{k}\sum\limits_{l=1}^{n-k}\{1-(1-d)^{l}\}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-k-l}\left(1-\frac{i}{N}\right)^{n-k-l}}{(n-k-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=n!\sum_{k=0}^{n-1}\frac{\left(\frac{(i-1)(1-d)}{N}\right)^{k}}{k!(n-k)!}\sum_{l=1}^{n-k}\binom{n-k}{l}\left[\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}-\left(\frac{1-d}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}\right]$
@@
$P_{d}(i;1st)$ $=e^{-(i-1)\frac{\mu Td}{N}}-e^{-i\frac{\mu Td}{N}}$
$=(e^{\frac{\mu Td}{N}}-1)e^{-i\frac{\mu Td}{N}}.$ (29)
The remaining probability
$1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td}$ (30)
is the probability of detecting no photon in a complete reference cycle. In that case, this particular reference cycle has no contribution to the random number. Therefore, the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
@@
Figure 5: Plots for the probabilities of detecting the first photon in $256$ time-bins given by Eq. (31). Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detection efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Note that $f(1)$ is the probability $P_{1}$ in Ref. (10.1063/1.4863224, , Eq. (4)). As their aim was to find the minimum entropy for deciding the amount of post-processing, they only derived an upper bound to $P_{1}$. However, we aim to reduce the amount of post-processing. Therefore, we have derived the complete probability mass function to see how the device parameters, $\mu,T$ and $d$ affect the distribution. We also discuss the effect of all the parameters separately in Section VII. Figure 5 shows how $\mu Td$ affects the probability (31). Here we consider $N=256$. From this plot, it is clear that the uniformity increases as $\mu Td$ decreases.
Although this uniformity can be increased by post-processing the raw data, as mentioned above, this reduces the random number generation rate and increases the possibility of side-channel attack. Since here we want to minimize the post-processing, the main goal in designing a QRNG is
$\underset{\mu,T,d}{minimize}\ g(\mu,T,d):=\mu Td,$ (32)
@@
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section VII.
Next, let us consider QRNG in Fig. 3(II). Here, the probability, that the waiting time ($t_{w}$) between the arrival of two consecutive photons lies in the $i$-th bin, is given by
$P((i-1)t_{l}<t_{w}<it_{l})$
$=P(t_{w}>(i-1)t_{l})-P(t_{w}>it_{l})$
$=e^{-(i-1)\mu t_{l}}-e^{-i\mu t_{l}},$ (33)
where $t_{l}$ is the length of each time-bin. If the detection efficiency is given by $d$, this probability is given by
@@
Since in the probability distribution of Eq. (VI.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (34) is identical with Eq. (VI.2). Therefore, the probability mass function corresponding to the probability Eq. (34) is also given by Eq. (31) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (12)). Since the distributions given by Eq. (31) are evaluated from the quantum properties of photons, this can be used in the goodness-of-fit test mentioned in Subsection V.3 to test the quantumness in the raw data generated by a QRNG.
@@
Figure 6: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution is $\epsilon$-random given by Eq. (36). Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
@@
Figure 6 is a visual representation of the Eq. (36) with $N=256$. It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to avoid post-processing while generating random numbers with $\epsilon$-randomness for a given $\epsilon$, one needs to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (36) has been satisfied.
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (31). These test results are provided in Section IX.
VII Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (32) of Subsection VI.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of
@@ VII Effect of $\mu,T,d$ on random number generation rate and cost
However, in the practical scenario, imperfections within used hardware components make ideal QRNG impossible. Therefore, the main goal in designing a QRNG is
[…]
@@
Figure 7: Total entropy (40) and minimum entropy (41) of the probability distribution (31). When the product of the expected number of photons, $\mu T$, within an external reference cycle $T$, and the detection efficiency $d$ is very small, $H(Q)$ and $H_{min}$ are very close to the entropy $H(U)=8$ of the uniform random distribution. Therefore, if $\mu Td$ is very small, the raw data generated by the QRNG would be close to the uniform random number.
@@
VII.1 Effect of the Expected Photon Numbers
@@
VII.1.1 Effect of the Exected Photon Number on Cost
@@
small value of μ high production $\text{small value of $\mu\implies$ high production cost}.$ (42)
VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (VI.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
@@ VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
Suppose, for some $\mu$, the expected number of external cycles to produce a random number of length $k$ bits, where each cycle contains $N$ time-bins, is given by $N_{c}$. Then $N_{c}=\langle X\rangle$, where $X$ is a random variable, whose probability distribution is negative binomial distribution, given by
@@
VII.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (32) of the Subsection VI.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
[…]
@@ VII.2 Effect of the timing error of the device on the External Reference Cycle
Now let us consider that $T$ is chosen to satisfy the relation $\frac{T}{N}\geq k\delta_{t}$, for some suitable $k$. Then the probability, $p_{e}$, that the detection happens in $i$-th bin but outside of the interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$ is bounded above as $p_{e}\leq\frac{2}{k}$. Here we have assumed uniform distribution as the time $\frac{T}{Nk}$ is very small. Therefore, for a fixed error tolerance, $p_{tol}$, $k$ can be chosen as $k_{tol}=\frac{2}{p_{tol}}$. In that case, the minimum reference cycle $T_{min}$ should be
@@
VII.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
@@
VII.4 Effect of Detection Efficiency
@@
VII.4.1 Effect of the Detection Efficiency on Cost
The value of $\mu Td$ can be decreased by choosing a detector with bad detection efficiency. It is very natural that if a detector can efficiently detect photons, its cost will also be high. Since the randomness increases if $\mu Td$ decreases, choosing small $d$ we can increase the randomness. This will also reduce the cost of the QRNG. Thus
small value of $\text{small value of }d\implies\begin{cases}\text{high randomness}\\ \text{low cost}\end{cases}$ (49)
VII.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
@@
Although a bad detector with a small detection efficiency, $d$, reduces production costs, this leads to a high chance of no photon detection in a particular external reference cycle. From Eq. (47) it can be easily seen that if $T$ decreases, the rate of generating random numbers will also decrease. The graph in Fig. 10 shows how this rate changes with $d$ for fixed $T$ and $\mu$. Therefore, it is not possible to choose a detector with arbitrarily small detection efficiency.
VIII Performance and Cost of the QRNG
We can see from Eq. (47) that random number generation rate $r\approx\mu d\log N$, where $\mu$ is the expected number of photons in unit time, $d$ is the efficiency of detector and $N$ is the number of time-bins. Figure 11 is a surface plot of Eq. (36) to show how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (32) of Subsection VI.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Also, from Eq. (42)and (49), we can write the cost of the QRNG as

### v6

@@ III.4 Experimental Results
IV Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
@@
The statistics discussed in Eq. (2) of Section II consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section V.
IV.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
@@
Currently available photon arrival time-based QRNGs quRNG ; TROPOS use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, would be $P(i;1st|n)=\left(1-\frac{i-1}{N}\right)^{n}-\left(1-\frac{i}{N}\right)^{n}$. This was observed by Nie et al. (10.1063/1.4863224, , Eq. (2)). However, we do not know the exact number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
@@
IV.2 Time-bin Statistics for Poissonian Photon with Imperfect Devices
@@
$P_{d}(i;1st)$ $=e^{-(i-1)\frac{\mu Td}{N}}-e^{-i\frac{\mu Td}{N}}=(e^{\frac{\mu Td}{N}}-1)e^{-i\frac{\mu Td}{N}}.$ (21)
The remaining probability $1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td}$ is the probability of detecting no photon in a complete reference cycle. In that case, this particular reference cycle has no contribution to the random number. Therefore, the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
@@
Figure 4: Plots for the probabilities of detecting the first photon in $256$ time-bins given by Eq. (22) with Python matplotlib library hunter2007matplotlib . Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detection efficiency. If all the optical components are ideal, the distribution should be given by the red dashed line.
Note that $f(1)$ is the probability $P_{1}$ in Ref. (10.1063/1.4863224, , Eq. (4)). As their aim was to find the minimum entropy for deciding the amount of post-processing, they only derived an upper bound to $P_{1}$. However, we aim to reduce the amount of post-processing. Therefore, we have derived the complete probability mass function to see how the device parameters, $\mu,T$ and $d$ affect the distribution. We also discuss the effect of all the parameters separately in Section V. Figure 4 shows how $\mu Td$ affects the probability (22). Here we consider $N=256$. From this plot, it is clear that the uniformity increases as $\mu Td$ decreases.
Although this uniformity can be increased by post-processing the raw data, as mentioned above, this reduces the random number generation rate and increases the possibility of a side-channel attack. Since here we want to minimize the post-processing, the main goal in designing a QRNG is
$\underset{\mu,T,d}{minimize}\ g(\mu,T,d)=\mu Td,$ (23)
@@
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section V.
Next, let us consider QRNG in Fig. 2(II). Here, the probability, that the waiting time ($t_{w}$) between the arrival of two consecutive photons lies in the $i$-th bin is given by $P((i-1)t_{l}<t_{w}<it_{l})=P(t_{w}>(i-1)t_{l})-P(t_{w}>it_{l})=e^{-(i-1)\mu t_{l}}-e^{-i\mu t_{l}}$, where $t_{l}$ is the length of each time-bin. If the detection efficiency is given by $d$, this probability is given by
@@
Since in the probability distribution of Eq. (21), $\frac{T}{N}$ is the length of the time-bins, Eq. (24) is identical with Eq. (21). Therefore, the probability mass function corresponding to the probability Eq. (24) is also given by Eq. (22) with $T=Nt_{i}$.
Therefore, these two types of QRNGs perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (2)). Since the distributions given by Eq. (22) are evaluated from the quantum properties of photons, this can be used in the goodness-of-fit test mentioned in Subsection III.3 to test the quantumness in the raw data generated by a QRNG.
@@
Figure 5: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution is $\epsilon$-random given by Eq. (26) with Python matplotlib library hunter2007matplotlib . Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
@@
Figure 5 is a visual representation of the Eq. (26) with $N=256$. It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to avoid post-processing while generating random numbers with $\epsilon$-randomness for a given $\epsilon$, one needs to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (26) is satisfied.
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (22). These test results are provided in Section VII.
V Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (23) of Subsection IV.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of
@@ V Effect of $\mu,T,d$ on random number generation rate and cost
However, in the practical scenario, imperfections within used hardware components make an ideal QRNG impossible. Therefore, the main goal in designing a QRNG is
$\textbf{Goal: }H(Q)\approx H(U).$ (28)
@@
Figure 6: Total entropy (30) and minimum entropy (31) of the probability distribution (22) with Python matplotlib library hunter2007matplotlib . When the product of the expected number of photons, $\mu T$, within an external reference cycle $T$, and the detection efficiency $d$ is very small, $H(Q)$ and $H_{min}$ are very close to the entropy $H(U)=8$ of the uniform random distribution. Therefore, if $\mu Td$ is very small, the raw data generated by the QRNG would be close to uniform random number.
@@
V.1 Effect of the Expected Photon Numbers
@@
V.1.1 Effect of the Expected Photon Number on Cost
@@
$\text{small value of $\mu\implies$ high production cost}.$ (32)
V.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (21), the probability, $P(0)$, of no photon detection within time $T$ is given by
@@ V.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
Suppose, for some $\mu$, the expected number of external cycles to produce a random number of length $k$ bits, where each cycle contains $N$ time-bins, is given by $N_{c}$. Then $N_{c}=\langle X\rangle$, where $X$ is a random variable, whose probability distribution is a negative binomial distribution, given by
@@
V.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (23) of the Subsection IV.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 8. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as (i1)(i-1)-th bin. Similarly, if the actual detection happens within time interval $i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
$\tau_{1}$$\tau_{2}$$\tau_{i}$$\tau_{N}$$\tau_{1}$$\tau_{2}$$\tau_{j}$$\tau_{N}$$\tau_{1}$$\tau_{2}$TTTTSignal Shifted Signal $\delta_{t}$
@@ V.2 Effect of the timing error of the device on the External Reference Cycle
Now let us consider that $T$ is chosen to satisfy the relation $\frac{T}{N}\geq k\delta_{t}$, for some suitable $k$. Then the probability, $p_{e}$, that the detection happens in $i$-th bin but outside of the interval $(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$ is bounded above as $p_{e}\leq\frac{2}{k}$. Here we have assumed uniform distribution as the time $\frac{T}{Nk}$ is very small. Therefore, for a fixed error tolerance, $p_{tol}$, $k$ can be chosen as $k_{tol}=\frac{2}{p_{tol}}$. In that case, the minimum reference cycle $T_{min}$ should be
@@
V.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
@@
V.4 Effect of Detection Efficiency
@@
V.4.1 Effect of the Detection Efficiency on Cost
The value of $\mu Td$ can be decreased by choosing a detector with a bad detection efficiency. It is very natural that if a detector can efficiently detect photons, its cost will also be high. Since the randomness increases if $\mu Td$ decreases, choosing small $d$ we can increase the randomness. This will also reduce the cost of the QRNG. Thus
$\text{small value of }d\implies\begin{cases}\text{high randomness}\\ \text{low cost.}\end{cases}$ (39)
V.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
@@
Although a bad detector with a small detection efficiency, $d$, reduces production costs, this leads to a high chance of no photon detection in a particular external reference cycle. From Eq. (37) it can be easily seen that if $T$ decreases, the rate of generating random numbers will also decrease. The graph in Fig. 9 shows how this rate changes with $d$ for fixed $T$ and $\mu$. Therefore, it is not possible to choose a detector with an arbitrarily small detection efficiency.
VI Simultaneous Analysis of Rate, Randomness, Entropy and Cost
We can see from Eq. (37) that random number generation rate $r\approx\mu d\log N$, where $\mu$ is the expected number of photons in unit time, $d$ is the efficiency of detector and $N$ is the number of time-bins. Figure 10 is a surface plot of Eq. (26) to show how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (23) of Subsection IV.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Also, from Eq. (32)and (39), we can write the cost of the QRNG as